Can you please help me find an equation of the line containing the given pair of points?
(1,3)and(3,4)
Thank you so very much!

Answer by kara!(17)

First find slope
the formula is:
m=(y₂-y₁ )⁄(x₂-x₁)
m= (4-3)/(3-1)
m=1/2
now use this formula to find an equation:
y-y₁ =m(x-x₁)
y-3=1/2(x-1)
y-3=1/2*x-1/2
y=1/2*x-5/2